A particle moves along the parabolic path Y equal to ax2ย in such a way that the X component of the velocity remains constant say c. The acceleration of the particle is
1. ackcap
2. 2ac2jcap
3.2ac2kcap
4.a2cjcap

Grade:11

1 Answers

Khimraj
5 years ago
accelerationย  = ax\hat{i}ย + ay\hat{j}
= (dvx/dt)\hat{i} + (dvy/dt)\hat{j}
ย = (dc/dt)\hat{i}ย + (d2y/dt2)\hat{j}
= 0 + d/dt[2ax(dx/dt)]\hat{j}
= d/dt[2acx]\hat{j}
= 2ac2ย \hat{j}
Hope it clears. If you like answer then please approve it.